OpenStudy (anonymous):

Find the area of the shaded sections. Click on the answer until the correct answer is showing.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is there a pic?

jimthompson5910 (jim_thompson5910):

use the formula A = (angle/360)*pi*r^2

jimthompson5910 (jim_thompson5910):

you'll have to find the area of each region separately, but they are congruent regions so you basically find the area of one region, then double that answer

OpenStudy (anonymous):

1/3(16(pi)

OpenStudy (anonymous):

16/3 (pi)

jimthompson5910 (jim_thompson5910):

A = (angle/360)*pi*r^2 A = (60/360)*pi*4^2 A = (1/6)*pi*16 A = (8/3)*pi That's the area of one region. Double it to get 2*(8/3)*pi = (16/3)*pi So you got it

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

and for this one?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

same idea, the radius is still 4 but now the angle is 90 degrees (instead of 60 degrees)

jimthompson5910 (jim_thompson5910):

oh wait, they want the triangles, not the pie-shaped pieces

OpenStudy (anonymous):

how did you find out it was 90?

jimthompson5910 (jim_thompson5910):

the area of one triangle is A = b*h/2 A = 4*4/2 A = 8

OpenStudy (anonymous):

ohhhh nvm ik how

jimthompson5910 (jim_thompson5910):

so both triangles combined make 2*8 = 16 square units

jimthompson5910 (jim_thompson5910):

yeah it shows a right angle (slightly covered up though)

OpenStudy (anonymous):

so the answer is 16?

OpenStudy (anonymous):

hwta happen to the other part?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

what about the part of 90/360 (pi)

OpenStudy (anonymous):

what about this one

jimthompson5910 (jim_thompson5910):

Area of Shaded Region = (Area of Larger Circle) - (Area of Smaller Circle)

OpenStudy (anonymous):

ohhh so like 16(pi)- 9(pi)

OpenStudy (anonymous):

Find the radius of a circle whose area equals the area of a rectangle that measures 2 ft. by 11 ft. (Use = 22/7)

OpenStudy (anonymous):

things are making much more sense to me noe

jimthompson5910 (jim_thompson5910):

more like 36pi - 9pi = 27pi

jimthompson5910 (jim_thompson5910):

but you have the right idea

jimthompson5910 (jim_thompson5910):

what's the area of the rectangle?

OpenStudy (anonymous):

thats what i meant to say

OpenStudy (anonymous):

Find the radius of a circle whose area equals the area of a rectangle that measures 2 ft. by 11 ft. (Use = 22/7)

OpenStudy (anonymous):

ummmm hold on idk how to get the h in the problem i never do knowhow

OpenStudy (anonymous):

h=height

jimthompson5910 (jim_thompson5910):

A = b*h A = 2*11 A = 22

jimthompson5910 (jim_thompson5910):

that's the area of the rectangle

jimthompson5910 (jim_thompson5910):

now use this area to find the radius of the circle

jimthompson5910 (jim_thompson5910):

this site is bugging out...

jimthompson5910 (jim_thompson5910):

is it working for you?

OpenStudy (anonymous):

the sie is working for me but im not sure how to get the correct answer i mean ik that (pi )(r)^2 equals the area but idk how to find the radius from the atinfo

OpenStudy (anonymous):

from that info *

jimthompson5910 (jim_thompson5910):

A = pi*r^2 22 = pi*r^2 solve for r

OpenStudy (anonymous):

sqrt7

jimthompson5910 (jim_thompson5910):

oh right and you use 22/7 for pi

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

22 = pi*r^2 22 = (22/7)*r^2 1 = (1/7)*r^2 7 = r^2 r = sqrt(7) you got it

OpenStudy (anonymous):

Find the difference in area between the circle and the triangle. Click on the answer until the correct answer is showing.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need to find the area of the circle and the triangle

OpenStudy (anonymous):

okay area of the circle is 9(pi)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

but how do i get h for the triangle idk how to find h for the problem A=1/2( b)(h)

jimthompson5910 (jim_thompson5910):

you would use trig to find the base and height of that triangle

OpenStudy (anonymous):

the base is 9 right ?

jimthompson5910 (jim_thompson5910):

cos(60) = adj/hyp cos(60) = adj/6 adj = 6*cos(60) adj = 3 So the shorter leg is 3 units

jimthompson5910 (jim_thompson5910):

the longer leg is 6*sin(60) = 6*sqrt(3)/2 = 3*sqrt(3)

jimthompson5910 (jim_thompson5910):

therefore, the area of the triangle is A = b*h/2 A = 3*(3*sqrt(3))/2 A = 9*sqrt(3)/2

OpenStudy (anonymous):

wouldnt it 9 (pi)-9/2 sqrt 3

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

that's the exact area

OpenStudy (anonymous):

okay one more i think

OpenStudy (anonymous):

Leave answer in exact form unless otherwise indicated. Find the area of the region between a regular hexagon with sides of 6" and its inscribed circle.

OpenStudy (anonymous):

how do i find the area of the hexagon?

jimthompson5910 (jim_thompson5910):

one sec, the drawing wasn't working out but I found a better drawing

OpenStudy (anonymous):

okie dokie :)

jimthompson5910 (jim_thompson5910):

ok here is a hexagon

jimthompson5910 (jim_thompson5910):

you can break it up into 6 pieces like so

jimthompson5910 (jim_thompson5910):

so each piece is a triangle with a base of 6 inches at this point, the height is unknown, but we can find it

jimthompson5910 (jim_thompson5910):

each interior angle is 180(6-2)/6 = 120 degrees. Split that in half to get 60 degrees So each triangle has 60 degrees in the lower portion like this